OpenStudy (anonymous):

I WILL MEDAL AND FAN PLEASE HELP! What is the slope of the line passing through the points (–3, 4) and (2, –1)?

OpenStudy (tkhunny):

Calculate the difference in the y values. Then what? Show YOUR work.

OpenStudy (anonymous):

Would the difference be -3?

OpenStudy (anonymous):

I won't do them for you, but I WILL tell you how to do them. Here is an explanation of how to do this type of problem. Reggie’s Math Mini-Lessons: Slope-intercept Question: How do I find the equation of a line given only two points on the line? How do I put this equation into slope-intercept form? Reggie Says: Okay. First, of all, let’s look at the formula for slope: y2 - y1 ______ = m x2-x1 This means that, for two points (x, y) , (x2, y2), you take the y coordinate from one point and subtract from the coordinate for the other point. Don’t worry if you get a negative number. Then take the x coordinates and do the same thing. Finally, take the number you got subtracting the y coordinates and put it over the number from the x coordinates, like a fraction. If I have two points, (5, 6) (4, 8) 8 - 6 = 2 2 _______ = ― 4 - 5 = -1 -1 So the slope is -2. Now, we can use the slope to find the y-intercept. Start at either point and count up or down the number of units indicated (a negative slope is down and to the right or up and to the left; a positive slope goes up-and-right OR down-and-left). Then count left or right the number of units indicated. Make sure you count TOWARD the y-axis. For a slope of -2, for example, count EITHER up 2 and left 1 or down 2 and right 1. Repeat this counting, in both directions if necessary, until you reach the y-axis. The point where you reach the y-axis is your y-intercept. Now we can put these into our slope-intercept formula: y = mx + b “m” is the slope you found with the first formula. “b” is the point where you touch the y-axis. Leave x and y alone. Special Rules: a slope of “zero” is a horizontal line. A line like “x = 6” is a line with an undefined slope. Guided Example: Find, in slope-intercept form, the line through (1, 5) and (3, 7) 7 - 5 = 2 3 - 1 = 2 2 / 2 = 1 So the slope is 1. Now, to find the y-intercept, I take one of the points (1, 5), which is the point closest to the y-axis, and count one space down and one space to the left (both points, in this case, are right of the y-axis). This puts me at (0, 4). So the y-intercept is 4. y = 1x + 4 y = x + 4 Now you try these (no peeking till you’ve done them!): Find, in slope-intercept form, the equation of the line passing through the points given a) (6, 4) , (5, 6) b) (7, 1), (5, 0) c) (8, 7), (7, 8) d) (0, 5), (2, 1) Answers: a) y = -2x + 16 b) y = 1/2 x + approx. .5 c) y = - x + 15 d) y = -2x + 5

OpenStudy (anonymous):

there it should tell you how to do it

OpenStudy (anonymous):

Because 4 - -1 would equal 3? Right? @tkhunny sorry @Devilwolf3000 But, you copied that off of somewhere else.

OpenStudy (anonymous):

well i need help i am stuck with this question

OpenStudy (anonymous):

Unit 3 | Lesson 6 Lesson Checkpoint: Show What You Know Question Navigator 4. Which equation can be used to find the answer? Joelle's stamp book has room for 54 stamps. She already has 38 stamps in the book. How many more stamps does she need to fill the book? A. s – 38 = 54 s = 92 92 stamps B. 2 · s = 54 s = 27 27 stamps C. s + 38 = 54 s = 16 16 stamps D. s ÷ 2 = 38 s = 76 76 stamps

OpenStudy (anonymous):

minus for a times for c division for d

OpenStudy (tkhunny):

4-(-1) = 4+1 = 5